OpenStudy (anonymous):

A volume of 26.65 ± 0.04 mL of HNO3 solution was required for complete reaction with 0.8328 ± 0.0007 g of Na2CO3, (FM 105.988 ± 0.001). Find the molarity of the HNO3 and its absolute uncertainty.

OpenStudy (aaronq):

Are you having problems finding the molarity of the uncertainty?

OpenStudy (anonymous):

Well I need both the molarity and the uncertainty and I am not sure how to do any of it.

OpenStudy (aaronq):

can you write an equation for the reaction?

OpenStudy (anonymous):

2H+ + Na2CO3 -> 2Na+ + H2O+ CO2

OpenStudy (aaronq):

so first were gonna find the volume, then we'll do the uncertainty after. find the moles of Na2CO3, then using the stoichiometric coefficients find the moles of HNO3

OpenStudy (aaronq):

\(n_{Na_2CO_3}=\dfrac{m_{Na_2CO_3}}{M_{Na_2CO_3}}\), btw.

OpenStudy (anonymous):

I got 0.007857 moles of Na2CO3. Then I couldn't find the stoichiometric coefficients. The one for Na2CO3 is 1mole but I dont know what the one is for HNO3?

OpenStudy (aaronq):

the stoichiometric coefficients are present in the balanced equation, now you wanna find moles of HNO3 (or as it's written in your equation H+). build a ratio, like this: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients (in red), \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\)

OpenStudy (anonymous):

so is the H+ the same as HNO3?

OpenStudy (aaronq):

for the purpose of the reaction, yes, because HNO3 is a strong acid (fully dissociates) and it's monoprotic.

OpenStudy (anonymous):

So now I have .015714 moles of HNO3 and ).02665L of HNO3.

OpenStudy (anonymous):

Which divid moles and L I get my molarity

OpenStudy (aaronq):

great, now find the molarity \(M=\dfrac{n_{solute}}{L_{solution}}\)

OpenStudy (anonymous):

now how do I get the absolute uncertainty?

OpenStudy (aaronq):

okay, use these equations:

OpenStudy (aaronq):

i'll show you the first calculation.

OpenStudy (aaronq):

m=0.8328 ± 0.0007 g ; M 105.988 ± 0.001 g/mol a \(\Delta a\) b \(\Delta b\) we divided, so we use \[s_y=n_{Na_2CO_3}*\sqrt{\dfrac{\Delta a}{a}+\dfrac{\Delta b}{b}}\]

OpenStudy (aaronq):

damn i forgot to write squared, it should be like this: \[s_y=n_{Na_2CO_3}*\sqrt{(\dfrac{0.0007 }{0.8328})^2+(\dfrac{0.001}{105.988}})^2\]

OpenStudy (aaronq):

does it make sense?

OpenStudy (anonymous):

I think so, I am trying it right now.

OpenStudy (anonymous):

so the n is the moles right?

OpenStudy (anonymous):

so where is the +or - 0.04

OpenStudy (aaronq):

yeah, thats why you have to do the calculation first, because you need the actual answers to compute the uncertainty. the 0.04 is used later when you're calculating the molarity

OpenStudy (aaronq):

These types of calculations can be a pain because you have to do them for every step an uncertainty is given

OpenStudy (anonymous):

so I got 6.6*10^-6 is that right?

OpenStudy (aaronq):

0.00000660465 yes

OpenStudy (anonymous):

It said it was wrong?

OpenStudy (anonymous):

the second part.

OpenStudy (aaronq):

which is the second part?

OpenStudy (anonymous):

the absolute uncertainty

OpenStudy (aaronq):

that's not the final answer, you have to do it again when you're finding molarity

OpenStudy (anonymous):

ok........

OpenStudy (aaronq):

use the uncertainty from: moles of Na2CO3 = \(0.00785 \pm6.6*10^-6 \) moles as the uncertainty in moles of HNO3, then use: volume of 26.65 ± 0.04 mL of HNO3 for the next calculation

OpenStudy (anonymous):

Ok I got it as 0.001 thanks!

OpenStudy (aaronq):

good stuff !